The length of one of the equal legs of an isosceles triangle is 8 cm less than 4 times the length of the base. If the perimeter is 29 cm, find the length of one of the equal legs.
Answer by Paul(988)

Let the "one side" be 4x-8
base = x
Equation:
2(4x-8)+x=29
8x-16+x=29
9x=45
x=5
4(5)-8=12
Hence, the base is 5cm and the length is 12cm
Paul.